A train on a straight track starts from rest t station A. It accelerates for 20 seconds with an acceleration of 5 m/s^2 and then continues to travel at the velocity reached for two minutes. The train then decelerates at 5 m/s^2 for 20 seconds before coming to a stop at the next station B. What is the distance between the two stations?
Answer #1

at station A train is at rest

for 20 sec

d1 = ut+1/2at^2

d1 = 1/2*5*20^2 = 1000 m

at the end of 20 s speed is v1 = a*t = 5*20 = 100 m/s

for next 20 min it continues with the speed of 100 m/s

d2 = 100*2*60 = 12000 m

for next 20 sec

d3 = 100*20-1/2*5*20^2 = 2000-1000 = 1000 m

and finally comes to rest at station B

so total distance between A and B is

d = d1+d2+d3

d = 1000+12000+1000 = 14000 m = 14 km